OpenStudy (anonymous):

integrate from 0 to 1 for [x/(x^2 +2)^2]

OpenStudy (mimi_x3):

\[\large \int\limits\frac{x}{(x^{2}+2)^{2}} \] It's a u- substitution. Let \(u = x^2+2\) => \(\frac{du}{2x} \)

sam (.sam.):

\[\frac{du}{2}=xdx\]

sam (.sam.):

Not \[\frac{du}{2x}\] :)

OpenStudy (mimi_x3):

Derivative of x^2 is 2x.. it is du/2x

sam (.sam.):